OpenStudy (anonymous):

What is the solution to the equation? 4.8 + x - 5.9 = 3.6

OpenStudy (anonymous):

This is not hard. You can do it. Bring all the number to the right side and you will get x.

OpenStudy (anonymous):

Bring all the numbers to the right. ._. So .. 4.8 - 5.9 = 3.6 ..Doesn't make sense. So..? I must be doing it wrong.

OpenStudy (anonymous):

Add 5.9 to the left and right: 4.8 + x - 5.9 + 5.9 = 3.6 + 5.9 Take away 4.8 from both side: 4.8 - 4.8 + x = 9.5 - 4.8 Simplify: x = 4.7 There you go :)

OpenStudy (anonymous):

Oh ..okay, I see. Thank you very much. ^^